I would like to know how to
SHOW THAT THIS GRAPH CAN ALSO BE REPRESENTED BY THE ARITHMETIC SEQUENCE EQUATION
an=a1+(n-1)d

Answer by greenestamps(13200)



I'm not sure what you mean by "the arithmetic sequence equation"....

a1 = a1
a2 = a1+(2-1)d = a1+d
a3 = a1+(3-1)d = a1+2d
a4 = a1+(4-1)d = a1+3d
...

There is a common difference of d between successive terms; that makes the sequence arithmetic....

If that doesn't answer your question, re-post it, making it clear what you are looking for.